# Cacophony: An Improved Contrastive Audio-Text Model

### v2

@@ V-C2 Experimental setup
We evaluate on Clotho-AQA [28] and Music-AQA, an audio-related subset of Music-AVQA dataset [55]. Clotho-AQA consists of 1,991 audio sounds randomly selected from Clotho, with six textual question-answer pairs related to each audio file. Four questions out of the six are designed to have ‘yes/no’ binary answers, while the remaining questions have other 828 single-word answers, which we treat as binary and multi-class classification respectively. The Music-AQA subset is composed of 6,319 sixty-second segments extracted from YouTube videos of musical performances. For each audio clip in the Music-AQA dataset, a single question is manually created to evaluate counting or comparative abilities, such as “Is the accordion playing longer than the saxophone?” and “How many musical instruments are audible in the video?” Similar to the Clotho-AQA dataset, the responses in the Music-AQA split are designed to provide binary answers or single-word numerical responses, ranging from ‘zero’ to ‘nine’.
We evaluate closed AQA through training small MLPs on top of frozen pretrained encoders. In particular, we first extract embeddings from the frozen audio and text encoders from the contrastive models, and then we concatenate them into one vector and pass it through a 4-layer MLP for binary or multi-class classification.
We also compare our approach to baseline model MWAFM [54] designed specifically for AQA that has achieved state-of-the-art performance on the Clotho-AQA and Music-AQA benchmarks. This model differs from contrastive models in two ways: firstly, its audio and text encoders are pretrained independently, and secondly, after extracting frame-level audio and token-level word embeddings, these embeddings are processed through fused attention layers, which is then followed by MLP layers. We follow the train, validation and test splits provided in the original Clotho-AQA and Music-AQA datasets. We use $R@k$ to evaluate the performance of the Clotho-AQA dataset because the size of the answer vocabulary is high (828). We use accuracy as a metric for MUSIC-AQA dataset because its single-word answer vocabulary consists of only 10 possible answers to different questions.
@@ V-C3 Result
In our evaluations, detailed in Table III, we benchmark fine-tuned contrastive models together with a state-of-the-art AQA baseline. Although Cacophony outperforms other contrastive models in binary question tasks, it performs less well in single-word classification, particularly when compared to the specialized AQA baseline. Cacophony uses pooled embeddings from a frozen text encoder, in contrast to the baseline, which learns to aggregate over granular text features  [54]. Given that the text in AQA task, i.e., a question, is out-of-distribution for our text encoder, the frozen pooled embeddings may not precisely capture the question’s semantics. A trainable pooler could offer improved pooled embeddings by leveraging the fine-grained output of the encoder.
@@ V-D1 Task definition
Audio classification is the process of categorizing different sound types on given audio recordings. To evaluate the generalization capability of the contrastive models on unseen datasets, we explore their zero-shot transfer ability on these audio classification tasks following previous practices [12, 16, 53]. This setup resembles audio-to-text retrieval task defined in Sec. II, with the key difference being the use of a predefined set of labels as text descriptions.
@@ V-D2 Experimental setup
We evaluate the models’ zero-shot classification accuracy using four datasets: VGGSound-test, TUT Acoustic Scenes [56], ESC-50 [49], and UrbanSound8K [50]. These datasets collectively cover a wide range of sound events. An overview of each dataset is provided in Tab. V. For each dataset during classification, we use the names of all classes within that dataset as the set of possible text descriptions, and as in audio-to-text retrieval, select the matching text based on cosine similarity. Since the text descriptions seen during training are relatively verbose compared to the often-single-word labels in classification tasks, we craft individual prompt templates for each dataset. For instance, in the context of sound event classification, we use a template such as “This is a sound of [label]”, while for acoustic scene classification, we use “This sound is on [label].” We use Top-1 accuracy as the evaluation metric.
TABLE V: Task descriptions for zero-shot classification and HEAR benchmark. Tasks with symbol † are used in zero-shot classification, tasks with ‡ are used in both tasks and the rest datasets are only used in HEAR benchmark.
Task name Task description Dur. (sec) # of Classes
VGGSound† Classification of sound events from YouTube 10.0 310
UrbanSound8K† Urban sound event classification 4.0 10
TUT Acoustic Scenes† Acoustic scenes classification 10.0 15
Speech Commands Spoken commands classification 1.0 12
NSynth Pitch Pitch classification of synthesized sounds 4.0 88
Beijing Opera Percussion Beijing Opera percussion instruments recognition 4.8 4
CREMA-D Speech emotion recognition 5.0 6
ESC-50‡ Environmental sound classification 5.0 10
Gunshot Triangulation Localization of microphone recording a gunshot 1.5 7
GTZAN Genre Music genre classification 30.0 10
GTZAN Music Speech Classification of audio into music or speech 30.0 2
LibriCount Multiclass speaker count identification 5.0 11
Vocal Imitations Match a vocal imitation to the imitated sound 11.3 302
VoxLingua107 Top10 Spoken language identification 18.6 10
MAESTRO 5h Event level music transcription 120 87
DCASE 2016 Task 2 Office sound event detection in synthesized scenes 120 11
@@ V-D3 Result
We compare our model, Cacophony, against all baseline models on zero-shot classification on a variety of audio classification benchmarks, presented in Table IV. Overall, Cacophony exhibits competitive performance for general sound categories on ESC-50, UrbanSound8K, and TUT-Acoustic Scenes. Since our model is trained on audio sampled at 16,000 Hz, training on full-bandwidth audio recordings could potentially improve its performance in zero-shot classification tasks, as demonstrated in previous work [2].
@@ V-E1 Task definition
The HEAR benchmark is designed to evaluate the effectiveness of general audio representations across various domains, including speech, music and general sound. The HEAR benchmark comprises two principal task types: (1) scene-based tasks, which involve classifying an entire audio clip, and (2) event-based tasks, which aim at identifying specific sound events over time, i.e., predicting the start time, end time, and label for each sound event. Detailed information about these tasks is listed in Tab. V.
@@ V-E2 Experimental setup
We follow the fine-tuning strategies presented in [30] using hear-eval-kit [^6] https://github.com/hearbenchmark/hear-eval-kit. Specifically, in both scene-based and event-based tasks, the audio encoders from the audio-text models are frozen and used as the input feature vector to a shallow downstream MLP classifier. All of the downstream tasks are trained with the same random seed and hyper-parameters. For scene-based tasks, the final layer is a softmax activation with Cross-Entropy (CE) loss, whereas in event-based tasks, the final layer is a sigmoid activation with binary CE loss. During evaluation following [30], scene-based tasks are measured with classification accuracy. Event-based tasks are evaluated with event onset F-measure (FMS), which correlates better with human perceptual than framewise classification accuracy [60].
@@ V-E3 Result
We conduct a thorough system comparison, evaluating our models against existing state-of-the-art contrastive models using the HEAR benchmark, shown in Fig. 5. WavCaps-HTSAT and our Cacophony models show relatively balanced performance across different datasets but no single model is superior across all tasks. Our proposed system demonstrates good performance across various tasks, outperforming or closely matching existing approaches. Notably, it achieves significantly better accuracy than other contrastive models in the pitch classification tasks defined in NSynth Pitch. However, our Cacophony model underperforms ‘WavCaps-HTSAT’ baseline in several speech-related datasets, including Speech Commands and LibriCount. We believe that incorporating more speech-specific data would improve performance on these speech-related tasks.
Interestingly, in a comparison between our first stage Cacophony model (listed as ‘AudioMAE’ in Fig. 5) and our second stage model (listed as ‘Cacophony’), i.e., before and after the contrastive-captioning training, we notice a significant drop in performance across various tasks. We hypothesize that this drop is caused by the difference between MAE and contrastive training objective. Intuitively, while the MAE encourages the audio encoder to represent information/features relevant to reconstructing adjacent spectrogram patches, the contrastive objective only encourages the audio encoder to represent information/features relevant to selecting a text pair. For example, our dataset may not contain text descriptions related to speaker emotion, so the contrastive learning may not explicitly incentivize the audio encoder to produce different representations for different emotions.
@@
TABLE VII: Automated audio captioning results on test sets of AudioCaps and Clotho. All compared baselines are from WavCaps [18], the evaluation results for the baselines are from the reference. HTSAT-BART-FT is fine-tuned specifically on AudioCaps, and we also used it for recaptioning weakly/un-labeled datasets.
@@ V-F Automated audio captioning
TABLE VIII: Representative examples of automatic captioning by Cacophony on Clotho dataset.
@@ V-F1 Task definition
Automated audio captioning involves generating a textual description, or caption, of an audio signal in free-form text, moving beyond the constraints of predefined class labels or tags. At inference, our model utilizes temperature sampling to generate captions, adjusting the probability distribution of the next word with a temperature parameter. We use a temperature of 0.1. The baselines apply beam search with a beam size of 3 for generating text; it’s possible that using beam search with our model would improve results.
@@ V-F2 Experimental setup
We evaluate on the AudioCaps and Clotho datasets. For evaluation metrics, we use the Microsoft COCO Caption Evaluation package [61], which includes BLEUn, ROUGEL, METEOR, CIDEr, SPICE, and SPIDEr.
@@ V-F3 Result
We benchmark our captioning head against ‘HTSAT-BART’ and ‘CNN14-BART’ [18, 62] which are pretrained on WavCaps. Both of these models have shown good performance in the automated audio captioning tasks on the Clotho and AudioCaps datasets. We also include the synthetic audio captioner for our collected Youtube data introduced in Sec. II, ‘HTSAT-BART-FT’ which has been additionally fine-tuned on AudioCaps. The results of this comparison are detailed in Table VII. In the AudioCaps dataset, our model exhibits performance that is competitive with the baseline models. However, there is a noticeable performance gap when compared to our “teacher model” HTSAT-BART-FT. On the Clotho dataset, Cacophony’s performance is notably lower than both CNN14-BART and HTSAT-BART. We have included a selection of sampled ground-truth captions alongside the generated captions that resulted in low metric scores in Table VIII. Upon reviewing these generated captions, we identify two primary causes for our model’s weak performance in reference-based metrics: First, the textual output from Cacophony diverges significantly from the distribution of the ground-truth text. These examples highlight a difference in writing style, with the generated text lacking the detailed descriptions characteristic of Clotho’s captions. It is expected that using a captioning model fine-tuned on Clotho to generate our pretraining data would enhance the downstream performance on this dataset, albeit potentially at the expense of performance on AudioCaps. Secondly, there is evidence that our audio encoder does not successfully detect all sound events, which contributes to the lower scores. For example, in the case of the middle example shown in Tab. VIII, the “birds chirping far away” sound event is not correctly identified by our model.
The results indicate that our model is capable of generating captions that closely approach the performance of the “teacher model”, but is also heavily influenced by the biases present in our collected pretraining dataset. This bias could potentially be mitigated by using synthetic captions that more closely resemble human-written captions.
Figure 6: UMAP visualization of text and audio embeddings on the validation set. (a) Evolution of modality gap during our second-stage training, $T$ stands for the temperature parameter in Eq. 1. (b) Visualization of modality gap of existing audio-text models.
@@ V-G Modality gap
Liang et al [63] define the modality gap as the difference between the centroids of audio and text embeddings:
$\Delta_{gap}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}-\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{y}_{i},$ (5)
where $\boldsymbol{x}_{i}$ and $\boldsymbol{y}_{i}$ denotes the L2-normalized audio and text embeddings for for the $i$-th sample in a given dataset. The magnitude of this gap provides a quantitative measure of the discrepancy between the audio and text embedding spaces. This gap is shown to arise from the general inductive bias of deep neural architectures, random initializations and training objective [63]. Although the contrastive training objective brings matched text-audio pairs closer than non-matched pairs, it does not necessarily bring audio and text into a joint multi-modal space globally. Another factor may be the noisy correspondence of audio-text pairs, as pointed out in Luong et al. [64]. The existence of the modality gap in contrastive image-text models can impact its transferability to downstream tasks [63]; we hypothesize that it may also exist in contrastive audio-text models. In our experiment, we track the evolution of the modality gap by plotting paired text and audio embeddings, as illustrated in Fig. 6(a). We observed that the modality gap begins to increase once the validation loss plateaus after 120k iterations. When evaluating the text-audio retrieval task on the validation split, we find that checkpoints with smaller modality gap do not necessarily guarantee a significant better performance. However, a model with a narrower gap may offer advantages in applications where embedding replacement across modalities is required, such as in text-to-audio generation [8, 9] or language-guided source separation [10, 11]. We also visualize the modality gaps of existing contrastive models, as shown in Fig. 6(b). In models like LAION-CLAP and WavCaps-HTSAT, we notice that text and audio embeddings form distinct clusters.
To summarize, our preliminary analysis on the modality gap shows that increases over the later stages of training and is present in contrastive audio-text models from previous works. We hypothesize that the modality gap could serve as guidance for selecting model checkpoints. We leave the question of the specific impact of modality gap on downstream performance to future work.
VI Ablation studies
TABLE IX: Ablations on text preprocessing, trained on openSFX, evaluated on test split of AudioCaps and Clotho.
Text pre-processing AudioCaps Clotho
@@ VI Ablation studies
TABLE X: Ablations on the use of dataset for different stage training, $V_{x}$-$V_{y}$ represents that $V_{x}$ dataset is used for 1st stage MAE training and $V_{y}$ dataset is used for 2nd stage captioning contrastive training.
Index Dataset AudioCaps Clotho
@@ VI-1 Text cleaning
In this section, we first explore text pre-processing methods for the contrastive-captioning training stage, using only the Freesound, AudioCaps, and Clotho datasets. To be specific, the compared text pre-processing methods include: (1) implementing LLM cleaning as detailed in Sec. II-B; (2) conducting recaptioning based solely on audio, as described in Sec. II-C; (3) directly using raw text within a predefined maximal text length. For recaptioning the Freesound dataset (2), we use CNN-14-BART from WavCaps [18], pretrained on WavCaps and then fine-tuned on Clotho, as its domain matches with Freesound
@@
The experimental result can be found in Tab IX. When compared to the raw text baseline, LLM cleaning generally improves performance for both text-to-audio and audio-to-text retrieval tasks on Clotho datasets, and achieves similar performance on AudioCaps dataset. Recaptioning yields mixed results; it improves performance on the Clotho dataset but slightly reduces accuracy on AudioCaps dataset.
@@
For both text preprocessing methods, improvements are more consistent and significant in Clotho than in AudioCaps. However, we believe that Clotho evaluation is a more reliable indicator of overall performance than AudioCaps, as the dataset used in this ablation aligns closely with Clotho and is out-of-distribution for AudioCaps. Though one might argue out-of-distribution evaluation is a useful analysis, we argue it is less relevant when (1) almost the entirety of the train dataset is from a particular source (in this case Freesound, of which Clotho is a subset) and (2) this is only an ablation experiment, and final experiments will include a much more diverse mixture of sources.
Overall, the LLM text-cleaning approach shows strong audio-text retrieval performance, with the best performance on AudioCaps and competitive with recaptioning on Clotho. Furthermore, LLM text-cleaning retains more details about the audio, and this may not be fully captured by the retrieval task defined on AudioCaps and Clotho datasets. We believe these observations support using LLM text cleaning for our final model, but we also hypothesize that combining cleaned descriptions with recaptioned descriptions (possibly again using a language model) may yield further improvements.
TABLE XII: Details on the different versions of dataset used for ablation experiments.
Dataset Clean Labeled Noisy-Labeled Weakly/Un-Labeled Duration (khour)
$V_{1}$ AudioCaps, Clotho, OpenSFX Freesound - 4.2
$V_{2}$ AudioCaps, Clotho, OpenSFX Freesound ACAV2M 7.8
$V_{3}$ AudioCaps, Clotho, OpenSFX Freesound AudioSet, ACAV2M 13.2
VI-2 Dataset scale
In our ablation study centered on different dataset decisions, we test datasets of different sizes, designated as $V_{1}$, $V_{2}$ and $V_{3}$, with each subsequent dataset being larger than its predecessor. Detailed information regarding the various dataset versions is provided in Table XII. In this section, we aim at investigating the influence of dataset scale in each training stage. The results for audio-text retrieval performance are presented in Table X.

### v3

@@ V-C2 Experimental setup
We evaluate our approach on two datasets: Clotho-AQA [29] and Music-AQA [53]. Clotho-AQA contains 1,991 audio samples, each with six questions. Four questions have yes/no answers, while two have single-word answers. Music-AQA, derived from the Music-AVQA dataset, includes 6,319 one-minute audio clips from musical performances. Each clip has one question assessing counting or comparison skills. Both datasets feature binary or single-word answers, with Music-AQA responses ranging from ‘zero’ to ‘nine’. We evaluate closed-ended AQA through training small MLPs on top of frozen pretrained encoders. In particular, we first extract embeddings from the frozen audio and text encoders from the contrastive models, and then we concatenate them into one vector and pass it through a 4-layer MLP for binary or multi-class classification.
In addition to several audio-text models, we also compare our approach to a baseline model [52] designed specifically for AQA that achieves state-of-the-art performance on the Clotho-AQA and Music-AQA benchmarks. This model differs from contrastive models in two ways: First, its audio and text encoders are pretrained independently, and second, after extracting frame-level audio and token-level word embeddings, these embeddings are processed through fused attention layers followed by MLP layers. We follow the train, validation and test splits provided in the original Clotho-AQA and Music-AQA datasets. We use $R@k$ to evaluate the performance on the Clotho-AQA dataset because the size of the answer vocabulary is large (828). We use accuracy to evaluate the performance on the Music-AQA dataset because its single-word answer vocabulary consists of only 10 possible answers to different questions.
@@ V-C3 Result
In our evaluations, detailed in Table III, we benchmark fine-tuned contrastive models together with a state-of-the-art AQA baseline. Although Cacophony outperforms other contrastive models in binary question tasks, it performs less well in single-word classification compared to the specialized AQA baseline. We believe this is because Cacophony uses pooled embeddings from a frozen text encoder, while the baseline learns to aggregate over granular text features  [52]. As the text in AQA task, i.e., a question, is out-of-distribution for our text encoder, the frozen pooled embeddings may not precisely capture the question’s semantics; A trainable pooler could offer improved pooled embeddings by leveraging the fine-grained output of the encoder.
@@ V-D1 Task definition
Audio classification is to categorize audio recordings into different sound types. To evaluate the generalization capability of the contrastive models on unseen datasets, we explore their zero-shot transfer ability on these audio classification tasks following previous practices [14, 17, 51]. This setup resembles the audio-to-text retrieval task defined in Sec. V-B, with the key difference being the use of a predefined set of labels as text descriptions.
@@ V-D2 Experimental setup
We evaluate the models’ zero-shot classification accuracy using four datasets: VGGSound-test, TUT Acoustic Scenes [54], ESC-50 [47], and UrbanSound8K [48]. These datasets collectively cover a wide range of sound events. For each dataset during classification, we use the names of all classes within that dataset as the set of possible text descriptions, and as in audio-to-text retrieval, select the matching text based on cosine similarity. Since the text descriptions seen during training are relatively verbose compared to the often-single-word labels in classification tasks, we craft individual prompt templates for each dataset. For instance, in the context of sound event classification, we use a template such as “This is a sound of [label]”, while for acoustic scene classification, we use “This sound is on [label].” We use Top-1 accuracy as the evaluation metric.
@@ V-D3 Result
We compare our model, Cacophony, against all baseline models on zero-shot classification on a variety of audio classification benchmarks, presented in Table IV. Overall, Cacophony exhibits competitive performance for general sound categories on ESC-50, UrbanSound8K, and TUT-Acoustic Scenes. However, it achieves a lower accuracy on the VGGSound dataset. This relatively lower performance on VGGSound, which includes 310 classes, suggests that its granularity may be too fine for the model to differentiate different classes. This challenge is likely due to the presence of inaccuracies within the re-captioned audio descriptions in the training data, potentially resulting from both the LLM cleaning and automatic captioning processes. Additionally, since our model is trained on audio sampled at 16,000 Hz, training on full-bandwidth audio recordings could potentially improve its performance in zero-shot classification tasks, as demonstrated in previous work [2].
@@ V-E1 Task definition
The HEAR benchmark is designed to evaluate the effectiveness of general audio representations across various domains, including speech, music and general sound. The HEAR benchmark comprises two principal task types: (1) scene-based tasks, which involve classifying an entire audio clip, and (2) event-based tasks, which aim at identifying specific sound events over time, i.e., predicting the start time, end time, and label for each sound event.
@@ V-E2 Experimental setup
We follow the fine-tuning strategies presented in [30] using hear-eval-kit [^7] https://github.com/hearbenchmark/hear-eval-kit. Specifically, in both scene-based and event-based tasks, the audio encoders from the audio-text models are frozen and used as the input feature vector to a shallow downstream MLP classifier. During evaluation following [30], scene-based tasks are measured with classification accuracy. Event-based tasks are evaluated with event onset F-measure, which correlates better with human perceptual than framewise classification accuracy [58].
@@ V-E3 Result
We evaluate our models against existing state-of-the-art contrastive models using the HEAR benchmark, shown in Fig. 5. Our proposed system demonstrates good performance across various tasks, outperforming or closely matching existing approaches. Notably, it achieves significantly better accuracy than other contrastive models in the pitch classification tasks defined in NSynth Pitch. WavCaps-HTSAT and our Cacophony models show relatively balanced performance across different datasets but no single model is superior across all tasks. Cacophony underperforms WavCaps-HTSAT baseline in several speech-related datasets, including Speech Commands and LibriCount. We believe that incorporating more speech-specific data would improve performance on these speech-related tasks.
Interestingly, in a comparison between Cacophony (stage 1) and Cacophony (stage 2), i.e., before and after the contrastive-captioning training, we notice a significant drop in performance across various tasks. This suggests that the contrastive training objective, which encourages the audio encoder to extract features relevant to selecting a text pair, may not align with other audio classification objectives.
For event-based tasks, we do not compare our model with existing contrastive models, because they tend to perform poorly without manually computing shifting-windowed embeddings, as they typically use temporal pooling in convolutional blocks or downsampling in Swin blocks [37]. Instead, we compare our model against top-performing individual models (i.e., excluding systems that ensemble multiple models) on the sound event detection tasks listed on the HEAR leaderboard [^8] HEAR leaderboard (https://hearbenchmark.com/hear-leaderboard.html). We exclude ensemble systems for a fair comparison on individual model performance.
@@
TABLE VI: Automated audio captioning results on test sets of AudioCaps and Clotho. All compared baselines are from WavCaps [19], the evaluation results for the baselines are from the reference. HTSAT-BART-FT is fine-tuned specifically on AudioCaps, and we also used it for recaptioning weakly/un-labeled datasets.
@@ V-F Automated audio captioning
TABLE VII: Representative examples of automatic captioning by Cacophony on Clotho dataset.
@@ V-F1 Task definition
Automated audio captioning involves generating a free-form textual description for an audio signal, moving beyond predefined class labels or tags.
@@ V-F2 Experimental setup
At inference, our model utilizes temperature sampling to generate captions at a temperature of 0.1. The baselines apply beam search with a beam size of 3 for generating text. We evaluate on the AudioCaps and Clotho datasets. For evaluation metrics, we use the Microsoft COCO Caption Evaluation package [59], which includes BLEUn, ROUGEL, METEOR, CIDEr, SPICE, and SPIDEr.
@@ V-F3 Result
We benchmark our captioning head against HTSAT-BART and CNN14-BART baselines proposed in [19] which are pretrained on WavCaps. Both of these models have shown good performance in the automated audio captioning tasks on the Clotho and AudioCaps datasets. We also include the synthetic audio captioner introduced in Sec. II, HTSAT-BART-FT, which has been additionally fine-tuned on AudioCaps.
The results of this comparison are detailed in Table VI. In the AudioCaps dataset, our model exhibits competitive performance with that of the baseline models. However, there is still a noticeable performance gap when compared to our “teacher model” HTSAT-BART-FT. On the Clotho dataset, Cacophony’s performance is notably lower than both CNN14-BART and HTSAT-BART. In Table VII, we include a few captiong examples of Cacophony that received low scores. Looking at these examples, we identify two primary causes for our model’s weak performance: First, the textual output from Cacophony seems to show a different writing style from the ground-truths, lacking some detailed descriptions that are characteristic of Clotho’s captions. It is expected that using a captioning model fine-tuned on Clotho to generate our pretraining data would enhance Cacophony’s performance on this dataset, albeit potentially at the expense of performance on AudioCaps. Second, Cacophony’s audio encoder does not successfully detect all sound events. This is evidenced in the middle example, where the “birds chirping far away” event is not correctly identified by our model.
Figure 6: UMAP visualization of text and audio embeddings on the validation set. (a) Evolution of modality gap during our second-stage training, $T$ stands for the temperature parameter in Eq. 1. (b) Visualization of modality gap of existing models.
@@ V-G Modality gap
Although the contrastive training objective brings matched text-audio pairs closer than non-matched pairs, it does not necessarily bring audio and text into a joint multi-modal space globally. Liang et al. [60] define the modality gap as the difference between the centroids of audio and text embeddings: $\Delta_{gap}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{x}_{i}-\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{y}_{i}$, where $\boldsymbol{x}_{i}$ and $\boldsymbol{y}_{i}$ denote the L2-normalized audio and text embeddings for the $i$-th sample in a given dataset. The magnitude of this gap provides a quantitative measure of the global discrepancy between the audio and text embeddings. This gap is shown to arise from the general inductive bias of deep neural architectures, random initialization and training objective [60]. Another factor may be the noisy correspondence of audio-text pairs, since correctly matched pairs guide the training, while mismatched pairs provide misleading supervision, where the pairs are incorrectly matched, as pointed out in Luong et al. [61], where the correctly matched pairs guide the training, the noisy pairs incorrectly supervise the training. As shown in contrastive image-text models, the existence of modality gap can impact its transferability to downstream tasks [60]; we hypothesize that it may also be the case for contrastive audio-text models. In our experiment, we track the evolution of the modality gap by plotting paired text and audio embeddings, as illustrated in Fig. 6a. We observe that the modality gap begins to increase once the validation loss plateaus after 120k iterations. When evaluating the text-audio retrieval task on the validation split, we find that checkpoints with smaller modality gap do not necessarily guarantee significantly better performance. However, a model with a narrower gap may offer advantages in applications where embedding replacement across modalities is required, such as in text-to-audio generation [8, 9] or language-guided source separation [10, 11]. We also visualize the modality gaps of existing contrastive models, as shown in Fig. 6b. In models like LAION-CLAP and WavCaps-HTSAT, we notice that text and audio embeddings form distinct clusters.
TABLE VIII: Ablations on text preprocessing, trained on openSFX, evaluated on test split of AudioCaps and Clotho.
[…]
@@
TABLE IX: Ablations on the use of dataset for different stage training, $V_{x}$-$V_{y}$ represents that $V_{x}$ dataset is used for 1st stage MAE training and $V_{y}$ dataset is used for 2nd stage captioning contrastive training.
[…]
@@
TABLE X: Detailed information of training datasets in the dataset scale ablation study.
Dataset Clean Labeled Noisy-Labeled Weakly/Un-Labeled Duration
$V_{1}$ AudioCaps, Clotho, OpenSFX Freesound - 4.2
$V_{2}$ AudioCaps, Clotho, OpenSFX Freesound ACAV2M 7.8
$V_{3}$ AudioCaps, Clotho, OpenSFX Freesound AudioSet, ACAV2M 13.2
@@ VI-A Text cleaning
We first explore text pre-processing methods for the contrastive-captioning training stage, using only the Freesound, AudioCaps, and Clotho datasets. To be specific, the compared text pre-processing methods include: (1) implementing LLM cleaning as detailed in Sec. II-B; (2) conducting recaptioning based solely on audio, as described in Sec. II-C; (3) directly using raw text within a predefined maximal text length. For recaptioning the Freesound dataset (2), we use CNN-14-BART from WavCaps [19], pretrained on WavCaps and then fine-tuned on Clotho, as its domain matches with Freesound.
@@
The experimental result can be found in Table VIII. When compared to the raw text baseline, LLM cleaning generally improves performance for both text-to-audio and audio-to-text retrieval tasks on Clotho datasets, and achieves similar performance on AudioCaps dataset. Recaptioning yields mixed results; it improves performance on the Clotho dataset but slightly reduces accuracy on AudioCaps dataset.
@@
For both text preprocessing methods, improvements are more consistent and significant in Clotho than in AudioCaps. We believe that Clotho evaluation is a more reliable indicator of the overall performance than AudioCaps evaluation, as the training dataset used in this ablation aligns closely with Clotho and is out-of-distribution for AudioCaps.
VI-B Dataset scale
We perform an ablation study on the effect of training dataset size on model performance in the audio-text retrieval tasks (see Table IX). To do so, we create three datasets designated as $V_{1}$, $V_{2}$ and $V_{3}$ with increasing sizes. Detailed information of these datasets is provided in Table X.